# X. Wang1,2,+    Siming Fu1,+    Qihan Huang1,2    Wanggui He1    Hao Jiang1,† 1

### v1

@@ Appendix I Subject Interpolation
Appendix J Limitations
There are certain limitations in MS-Diffusion. The box-based indication of positions lacks precision, making it challenging to work effectively when the interaction between subjects is stronger. Moreover, the model requires explicit layout input during inference, and generating complex scenes becomes difficult. Though MS-Diffusion beats SOTA personalized diffusion methods in both single-subject and multi-subject generation, it still suffers from the influence of background in subject images. In the future, we will tackle the above limitations and extend the model to more applications.
Appendix K Societal Impact

### v2

@@ Appendix I Multi-subject Interaction
Appendix M Limitations
Figure 11: Qualitative examples when applying pseudo layout guidance during the inference. In this figure, pse-lg and pse respectively indicate whether layout prior is used.
There are certain limitations in MS-Diffusion. The box-based indication of positions lacks precision, making it challenging to work effectively when the interaction between subjects is stronger. Moreover, the model requires explicit layout input during inference, and generating complex scenes becomes difficult. Though MS-Diffusion beats SOTA personalized diffusion methods in both single-subject and multi-subject generation, it still suffers from the influence of background in subject images.
We explore a solution for explicit layout needs during inference. MS-Diffusion supports using the text cross-attention maps as the pseudo layout guidance. Specifically, as indicated in Figure 10, since the text cross-attention maps can reflect the area of each text token, we can replace the layout prior with them during the inference. In practice, we set a threshold to extract masks from text cross-attention maps and apply them after $T$ denoising steps. Before $T$, we experiment with completely disabling layout guidance or using a rough box as a layout prior. The results are presented in Figure 11. Although disabling layout guidance experiences a decline in subject consistency, it still demonstrates that explicit layout guidance during inference can be optimized. One direction for exploration is to enable the model to learn the layout during training.
Appendix N Societal Impact